# Unlearning regularization for Boltzmann Machines

### v2

@@ Abstract
Boltzmann Machines (BMs) are graphical models with interconnected binary units, employed for the unsupervised modeling of data distributions. When trained on real data, BMs show the tendency to behave like critical systems, displaying a high susceptibility of the model under a small rescaling of the inferred parameters. This behaviour is not convenient for the purpose of generating data, because it slows down the sampling process, and induces the model to overfit the training-data. In this study, we introduce a regularization method for BMs to improve the robustness of the model under rescaling of the parameters. The new technique shares formal similarities with the unlearning algorithm, an iterative procedure used to improve memory associativity in Hopfield-like neural networks. We test our unlearning regularization on synthetic data generated by two simple models, the Curie-Weiss ferromagnetic model and the Sherrington-Kirkpatrick spin glass model, and we show that it outperforms $L_{p}$-norm schemes. Finally, we discuss the role of parameter initialization.
@@ I Introduction
One limitation consists in the incompatibility between the true unknown distribution of the training data and the energy-based distribution that is employed to fit them. As a consequence, the training process cannot converge to the best possible generative performance, and only selects a good approximate model among many possible similarly good choices. Hence, regularization can be used to push the training process to yield specific desidered properties, useful for particular applications of the data analysis. Standard schemes for regularization include $L_{2}$- and $L_{1}$-norm penalties imposed on the pairwise interactions. However, these regularization schemes have a tendency to decrease the values of the interactions, and introduce systematic biases in the model.
@@
The goal of this study, motivated by the practical issues described above, is to propose a new regularization technique for BM learning that explicitly aims at increasing the robustness of the model under rescaling of its parameters, i.e. the coupling matrix $\bm{J}=\{J_{ij}\}$ and the local fields $\bm{h}=\{h_{i}\}$. This new tool, that we named unlearning regularization, appears to be more effective than other techniques, approaching a higher robustness performance of the neural network. The performance of the unlearning regularization is evaluated from artificial data generated by both a Curie-Weiss (CW) and a Sherrington-Kirkpatrick (SK) models with a small number of variables. In this way the original parameters are known, the fixed points of the learning equations are reached precisely and the useful quantities can be computed exactly.
@@
The structure of the article is the following. We first introduce generative modeling and Boltzmann Machines in section II. Some extensively employed regularization techniques (i.e. the $L_{p}$ methods) are then described in section II.2. Section III briefly describes the HU algorithm and its traditional use in associative memory models. The new unlearning regularization is then defined in section III.2. Its performance is analyzed for two different data sources in section IV: a CW model (section IV.1) and a SK model (section IV.2). In section IV.4 we compare the new regularization method to the standard $L_{p}$ regularization method, showing that it substantially improves the robustness under rescaling of the parameters while preserving a high similarity to the original model. The analysis is performed both at convergence of the regularization algorithm and in the condition of highest similarity with the ground-truth model, obtained via early-stopping in the training. The study of a particular limit of the regularization technique leading to HU follows in section V, which provides some insight on its behavior as a generic inference tool.
@@ II.1 Boltzmann Machines
where $J_{ij}$ are symmetric couplings and $h_{i}$ are the fields acting on each neuron site. We know, from statistical mechanics, that such a system at equilibrium at a temperature $\beta^{-1}$ obeys the following Gibbs-Boltzmann joint probability distribution
@@ III.1 Unlearning algorithm
We will now describe an unsupervised algorithm employed in the realm of associative memory models that will be our main focus in the following. Inspired by the brain functioning during REM sleep [26], the Hebbian Unlearning algorithm (HU) [8, 26, 9, 27, 10, 28, 29] is a training procedure for simple associative models. Consider a neural network having the same energy defined in Eq. (2). We assume the external local fields to be null, i.e. $h_{i}=0$, $\forall i$. The associative memory task consists in retrieving a set of random neural states $\{\vec{\xi}^{\mu}\}_{\mu=1}^{\alpha N}$, called patterns, through a zero temperature Monte Carlo dynamics initialized on some corrupted versions of them. The control parameter $\alpha$ is the load of the model. Associativity is improved by achieving larger basins of attraction of the patterns.
@@
    Run the asynchronous dynamics until convergence to a stable fixed point $\vec{S}^{*}$.
@@ IV Results
In order to benchmark our proposed regularization scheme, we consider a small sized network, i.e. $N\sim 18\div 20$, which allows us to consider the limit of infinite number of generated data, i.e. $M\to\infty$, and compute exactly the moments $\langle\cdot\rangle_{data}$. We can then reach the fixed points of Eqs. (26) and (27) precisely, whenever such points are admitted, with no errors due to the finite sampling. We can then compute the observables $C_{v}$ or $sD_{KL}$ more easily, in order to determine the accuracy and robustness of the inference. It must be underlined that our purpose in this framework is not to maximize the generalization properties of the model, because this goal would be trivially reached by the standard BM learning: it is rather to find the model that achieves the best compromise between accuracy and robustness.
@@
where we will mainly deal with $\beta=a$ and $\beta=1$. The inverse temperature of the data generating model is $\beta=\beta_{d}$. For given $\beta_{d}$, we can compute $c_{ij}^{d}$ exactly in the limit $M\to\infty$ by exact enumeration of the model configurations. We consider two kind of data generating models: a mean field fully connected ferromagnetic Ising network (i.e. the Curie-Weiss model) and a fully disordered spin glass network (i.e. the Sherrington-Kirkpatrick model [32]), which both display a critical point in the thermodynamic limit, manifested by a singularity of the specific heat at a given critical temperature. This singularity is smoothed in finite size systems, thus appearing as a peak in specific heat. In our small systems, solving the BM learning equations leads to the exact parameters of the generating model. Instead, solving Eqs. (26) and (27) leads to a non-trivial network with its own inferential performance in terms of accuracy and robustness.
@@ IV.2 Sherrington-Kirkpatrick model: training
We consider a network of $N=18$ spins with a data generation temperature $\beta_{d}=0.4$. For clarity of the results, we will use only one realisation of the parameters for the SK, having a peak in the specific heat at a $\beta$ slightly larger than $\beta_{d}$. The parameters are initialised as $J_{ij}(0)=c^{d}_{ij}$ and $h_{i}(0)=0$ $\forall i$.
@@
Figure 3: (a) Symmetric KL divergence (in log scale), as a function of time $t\lambda$, between the inferred Gibbs-Boltzmann distribution at unitary temperature and the one describing a SK model at $\beta_{d}=0.4$. (b) Slope $\omega$ obtained from the linear fit of the elements of $\bm{c}_{d}$ versus the elements of $\bm{c}_{1}$, as a function of time $t$. (c) Pearson coefficient $\rho$ between the correlation matrix $\bm{c}_{1}$ and the data correlation matrix $\bm{c}_{d}$ as a function of time $t$. The full line reports the case of Hebbian initialization of the couplings; the dotted line corresponds to a random initialization; the dashed line represents the initialization to $\bm{J}^{(0)}=0$. Choice of the parameters: $N=18$, $a=0.3$, $\lambda=0.06$.
Figure 4: (a) Symmetric KL divergence (in log scale), as a function of time $t\lambda$, between the Gibbs-Boltzmann distribution at unitary temperature, inferred via $L_{1}$ and $L_{2}$ regularizations, and the one describing a SK model at $\beta_{d}=0.4$. (b) Slope $\omega$ obtained from the linear fit of the elements of $\bm{c}_{d}$ versus the elements of $\bm{c}_{1}$, as a function of time $t\lambda$. (c) Pearson coefficient $\rho$ between the correlation matrix $\bm{c}_{1}$ and the data correlation matrix $\bm{c}_{d}$ as a function of time $t\lambda$. The full line reports the case of Hebbian initialization of the couplings; the dotted line corresponds to a random initialization; the dashed line represents the initialization to $\bm{J}^{(0)}=0$. Choice of the parameters: $N=18$, $\beta_{d}=0.4$, $\lambda_{1}=\lambda_{2}=0.001$, $\gamma_{1}=0.15$, $\gamma_{2}=1$.
@@
To compare the correlation matrices $\bm{c}_{1}$ and $\bm{c}_{d}$, a non-weighted fit of the elements of $c_{d}$ versus the corresponding elements of $c_{1}$ provides the slope $\omega(\bm{c}_{1},\bm{c}_{d})$, while the Pearson coefficient $\rho(\bm{c}_{1},\bm{c}_{d})$ measures the degree of correlation between the two matrices. A high accuracy is reached when both $\rho$ and $\omega$ are close to unity. Figures 3b and 3c report the evolution of $\omega$ and $\rho$ during training. As one can notice, there is a clear correspondence between the local minimum of the symmetric KL divergence and the point where $\omega=1$ and $\rho$ displays a local maximum. It is therefore interesting to consider stopping the training at this point (defined in practice by $\omega=1$), in order to achieve the best accuracy. We note that this is only possible with a Hebbian initialization of the parameters; while a null inizialization never reaches $\omega=1$, a random one reaches it but with a much lower Pearson coefficient $\rho$, resulting in poor accuracy. Note that the correspondence between the local minimum of $sD_{KL}$ and the best match of $\bm{c}_{d}$ and $\bm{c}_{1}$ is not valid for the entire range of $a$. Specifically, for the rest of our analysis we will avoid values of $a$ that are close to unity, because for this value the correspondence does not hold and we cannot define a good early-stopping criterion.
@@ IV.3 SK model: results at convergence
The couplings are optimized until convergence of the training for the three types of regularizations. Three quantities are measured after training: the symmetric Kullback-Leibler divergence $sD_{KL}(1,d)$ between the model at unitary temperature and the original SK model, the specific heat at unitary temperature $C_{v}(1)$ and the slope $\omega$ defined as in section IV.2. One should keep in mind that $\omega=1$ is reached only by the trivial choice of the regularization parameters $a=1$ and $\gamma_{1}=0$, $\gamma_{2}=0$; any other choice of these parameters leads to a reduced accuracy, by definition of regularization. Measures are repeated for different choices of the regularization parameters $\{a,\gamma_{1},\gamma_{2}\}$ and results are reported in fig. 5. Specifically, fig. 5a is a colour plot showing the performance of the network in the plane $(sD_{KL},C_{v}(1))$, where each point is coloured according to the slope $\omega$. The three lines correspond to the three regularizations: the unlearning line is generally closer to the origin than the $L_{1}$ and $L_{2}$ ones, suggesting a better compromise between accuracy and robustness. All the lines converge to the ground-truth when $sD_{KL}=0$, corresponding to absence of regularization ($a=1$ or $\gamma_{1}=0$ or $\gamma_{2}=0$).
@@ IV.4 SK model with early-stopping
Figure 7: Comparison at early-stopping of the performance of the unlearning, $L_{1}$ and $L_{2}$ regularization schemes. Each point has two coordinates: the symmetric Kullback-Leibler divergence $sD_{KL}$ and specific heat at unitary temperature $C_{v}(1)$. The colorbar relates each point to the standard deviation of the couplings $\sigma_{J}$. Arrows display the increase in the regularization rates of the algorithms. The dashed line indicates the specific heat of the original SK model that generated the data. Choice of the parameters: $N=18$, $\beta_{d}=0.4$, $a\in[0,0.63]\cup[1.9,3]$, $\gamma_{1}\in[0.5,0.4]$, $\gamma_{2}\in[0.3,4]$, $\lambda=0.01$, $\lambda_{1}=\lambda_{2}=0.001$.

### v3

@@ Abstract
Boltzmann Machines (BMs) are graphical models with interconnected binary units, employed for the unsupervised modeling of data distributions. When trained on real data, BMs show the tendency to behave like critical systems, displaying a high susceptibility of the model under a small rescaling of the inferred parameters. This behaviour is not convenient for the purpose of generating data, because it slows down the sampling process, and induces the model to overfit the training-data. In this study, we introduce a regularization method for BMs to improve the robustness of the model under rescaling of the parameters. The new technique shares formal similarities with the unlearning algorithm, an iterative procedure used to improve memory associativity in Hopfield-like neural networks. We test our unlearning regularization on synthetic data generated by two simple models, the Curie-Weiss ferromagnetic model and the Sherrington-Kirkpatrick spin glass model. We show that it outperforms $L_{p}$-norm schemes and discuss the role of parameter initialization. Eventually, the method is applied to learn the activity of real neuronal cells, confirming its efficacy at shifting the inferred model away from criticality and coming out as a powerful candidate for actual scientific implementations.
@@ I Introduction
While in theory BMs (with a sufficient number of hidden units) are universal approximators, in practical applications one is limited by the incompatibility between the true unknown distribution of the training data and the energy-based distribution that is employed to fit them. As a consequence, the training process cannot converge to the best possible generative performance, and only selects a good approximate model among many possible similarly good choices [5]. Hence, regularization can be used to push the training process to yield specific desidered properties, useful for particular applications of the data analysis. Standard schemes for regularization include $L_{2}$- and $L_{1}$-norm penalties imposed on the pairwise interactions. However, these regularization schemes have a tendency to decrease the values of the interactions, and introduce systematic biases in the model.
@@
The goal of this study, motivated by the practical issues described above, is to propose a new regularization technique for BM learning that explicitly aims at increasing the robustness of the model under rescaling of its parameters, i.e. the coupling matrix $\bm{J}=\{J_{ij}\}$ and the local fields $\bm{h}=\{h_{i}\}$. This new tool, that we named unlearning regularization, appears to be more effective than other techniques, approaching a higher robustness performance of the neural network. The performance of the unlearning regularization is evaluated from artificial data generated by both a Curie-Weiss (CW) and a Sherrington-Kirkpatrick (SK) models. We initially consider a small number of variables: in this way the original parameters are known, the fixed points of the learning equations are reached precisely and the useful quantities can be computed exactly. In addition, higher-dimensional and more realistic cases are also considered to validate the results.
@@
The structure of the article is the following. We first introduce generative modeling and Boltzmann Machines in section II. Some extensively employed regularization techniques (i.e. the $L_{p}$ methods) are then described in section II.2. Section III briefly describes the HU algorithm and its traditional use in associative memory models. The new unlearning regularization is then defined in section III.2. Its performance is analyzed for two different data sources in section IV: a CW model (section IV.1) and a SK model (sections IV.2, IV.3, IV.4 and IV.5). In sections IV.3, IV.4 and IV.5 we compare the new regularization method to the standard $L_{p}$ regularization scheme, showing that it substantially improves the robustness under rescaling of the parameters while preserving a high similarity to the original model. The analysis is performed both at convergence of the regularization algorithm and in the condition of highest similarity with the ground-truth model, obtained via early-stopping in the training, either in the small and large $N$ cases. The study of a particular limit of the regularization technique leading to HU follows in section V, which provides some insight on its behavior as a generic inference tool. Eventually, the unlearning method is tested on real biological data in section VI, confirming in practice the validity of our new regularization prescription.
@@ II.1 Boltzmann Machines
where $J_{ij}$ are symmetric couplings and $h_{i}$ are the fields acting on each neuron site. Even if several types of architectures are collected under the title of BMs, we will stick to the simplest model, the one endowed exclusively with visible input neurons that interact reciprocally and are subjected to external fields. In fact, our goal will be analyzing the evolution of the interactions between real neurons, which are not necessarily present in other types of machines.
We know, from statistical mechanics, that such a system at equilibrium at a temperature $\beta^{-1}$ obeys the following Gibbs-Boltzmann joint probability distribution
@@ III.1 Unlearning algorithm
We will now describe an unsupervised algorithm employed in the realm of associative memory models that will be our main focus in the following. Inspired by the brain functioning during REM sleep [27], the Hebbian Unlearning algorithm (HU) [9, 27, 10, 28, 11, 29, 30] is a training procedure for simple associative models. Consider a neural network having the same energy defined in Eq. (2). We assume the external local fields to be null, i.e. $h_{i}=0$, $\forall i$. The associative memory task consists in retrieving a set of random neural states $\{\vec{\xi}^{\mu}\}_{\mu=1}^{\alpha N}$, called patterns, through a zero temperature Monte Carlo dynamics [31] initialized on some corrupted versions of them. Such a rule for the dynamics reads
$S_{i}^{(t+1)}=\text{sign}\left(\sum_{j}J_{ij}S_{j}^{(t)}\right).$ (22)
The control parameter $\alpha$ is the load of the model. Associativity is improved by achieving larger basins of attraction of the patterns.
@@
    Run the zero temperature Monte Carlo dynamics, updating a randomly picked neuron at each time step according to Eq. (22), until convergence to a stable fixed point $\vec{S}^{*}$.
@@ IV Results
In order to benchmark our proposed regularization scheme, we first consider a small sized network, i.e. $N=\{18,20\}$, which allows us to consider the limit of infinite number of generated data, i.e. $M\to\infty$, and compute exactly the moments $\langle\cdot\rangle_{data}$. We can then reach the fixed points of Eqs. (27) and (28) precisely, whenever such points are admitted, with no errors due to the finite sampling. We can then compute the observables $C_{v}$ or $sD_{KL}$ more easily, in order to determine the accuracy and robustness of the inference. It must be underlined that our purpose in this framework is not to maximize the generalization properties of the model, because this goal would be trivially reached by the standard BM learning: it is rather to find the model that achieves the best compromise between accuracy and robustness. In the second instance, we generalize our analysis to the case of larger $N$, where quantities cannot be computed exactly.
@@
where we will mainly deal with $\beta=a$ and $\beta=1$. The inverse temperature of the data generating model is $\beta=\beta_{d}$. For given $\beta_{d}$, we can compute $c_{ij}^{d}$ exactly in the limit $M\to\infty$ by exact enumeration of the model configurations. We consider two kind of data generating models: a mean field fully connected ferromagnetic Ising network (i.e. the Curie-Weiss model) and a fully disordered spin glass network (i.e. the Sherrington-Kirkpatrick model [34]), which both display a critical point in the thermodynamic limit, manifested by a singularity of the specific heat at a given critical temperature. This singularity is smoothed in finite size systems, thus appearing as a peak in specific heat. For both models we will choose $\beta_{d}$ to be smaller than the real position of the peak of $C_{v}(\beta)$ because sampling is more efficient in the paramagnetic phase.
@@ IV.2 Sherrington-Kirkpatrick model: training
We consider a network of $N=18$ spins with a data generation temperature $\beta_{d}=0.4$. For clarity of the results, we will use only one realisation of the parameters for the SK, having a peak in the specific heat at a $\beta$ slightly larger than $\beta_{d}$. The parameters are initialised as $\bm{J}^{(0)}=\bm{c}_{d}$ and $\bm{h}^{(0)}=\bm{0}$ $\forall i$.
@@
Figure 3: (a) Symmetric KL divergence (in log scale), as a function of time $t\lambda$, between the inferred Gibbs-Boltzmann distribution at unitary temperature and the one describing a SK model at $\beta_{d}=0.4$. (b) Slope $\omega$ obtained from the linear fit of the elements of $\bm{c}_{d}$ versus the elements of $\bm{c}_{1}$, as a function of time $t$. (c) Pearson coefficient $\rho$ between the correlation matrix $\bm{c}_{1}$ and the data correlation matrix $\bm{c}_{d}$ as a function of time $t$. The full line reports the case of Hebbian initialization of the couplings; the dotted line corresponds to a random initialization; the dashed line represents the initialization to $\bm{J}^{(0)}=0$. The gray vertical line signs the location of the minimum of $sD_{KL}$ for the Hebbian inizialization. Choice of the parameters: $N=18$, $a=0.3$, $\lambda=0.06$.
Figure 4: (a) Symmetric KL divergence (in log scale), as a function of time $t\lambda$, between the Gibbs-Boltzmann distribution at unitary temperature, inferred via $L_{1}$ and $L_{2}$ regularizations, and the one describing a SK model at $\beta_{d}=0.4$. (b) Slope $\omega$ obtained from the linear fit of the elements of $\bm{c}_{d}$ versus the elements of $\bm{c}_{1}$, as a function of time $t\lambda$. (c) Pearson coefficient $\rho$ between the correlation matrix $\bm{c}_{1}$ and the data correlation matrix $\bm{c}_{d}$ as a function of time $t\lambda$. The full line reports the case of Hebbian initialization of the couplings; the dotted line corresponds to a random initialization; the dashed line represents the initialization to $\bm{J}^{(0)}=0$. Light blue and red vertical lines sign the locations of the minima of $sD_{KL}$ for the Hebbian initialization, respectively in the $L_{1}$ and $L_{2}$ methods. Choice of the parameters: $N=18$, $\beta_{d}=0.4$, $\lambda_{1}=\lambda_{2}=0.001$, $\gamma_{1}=0.15$, $\gamma_{2}=1$.
@@
To compare the correlation matrices $\bm{c}_{1}$ and $\bm{c}_{d}$, a non-weighted fit of the elements of $c_{d}$ versus the corresponding elements of $c_{1}$ provides the slope $\omega(\bm{c}_{1},\bm{c}_{d})$, while the Pearson coefficient $\rho(\bm{c}_{1},\bm{c}_{d})$ measures the degree of correlation between the two matrices. A high accuracy is reached when both $\rho$ and $\omega$ are close to unity. Note that the estimation of $\omega$ and $\rho$ carry the same computational cost, in fact at each algorithm time step we have
$\small\omega^{(t)}(\bm{c}_{1},\bm{c}_{d})=\frac{\rho^{(t)}(\bm{c}_{1},\bm{c}_{d})-\mu_{d}\cdot\mu_{1}(t)}{s_{1}^{2}(t)-\mu_{1}(t)^{2}},$ (45)
where $\mu_{d}$ is the average over the off-diagonal elements of $\bm{c}_{d}$ and $\mu_{1}(t),s_{1}^{2}(t)$ are, respectively, the average of the off-diagonal elements and the average of the off-diagonal squared elements of $\bm{c}_{1}$ at time $t$.
Figures 3b and 3c report the evolution of $\omega$ and $\rho$ during training. As one can notice, there is a clear correspondence between the local minimum of the symmetric KL divergence and the point where $\omega=1$ and $\rho$ approaches a local maximum. It is therefore interesting to consider stopping the training at this point (defined in practice by $\omega=1$), in order to achieve the best accuracy. We note that this is only possible with a Hebbian initialization of the parameters; while a null inizialization never reaches $\omega=1$, a random one reaches it but with a much lower Pearson coefficient $\rho$, resulting in poor accuracy. Note that the correspondence between the local minimum of $sD_{KL}$ and the best match of $\bm{c}_{d}$ and $\bm{c}_{1}$ is not valid for the entire range of $a$. Specifically, for the rest of our analysis we will avoid values of $a$ that are close to unity, because for this value the correspondence does not hold and we cannot define a good early-stopping criterion.
@@ IV.3 SK model: results at convergence
The couplings are optimized until convergence of the training for the three types of regularizations. Three quantities are measured after training: the symmetric Kullback-Leibler divergence $sD_{KL}(1,d)$ between the model at unitary temperature and the original SK model, the specific heat at unitary temperature $C_{v}(1)$ and the slope $\omega$ defined as in eq. 45. One should keep in mind that $\omega=1$ is reached only by the trivial choice of the regularization parameters $a=1$ and $\gamma_{1}=0$, $\gamma_{2}=0$; any other choice of these parameters leads to a reduced accuracy, by definition of regularization. Measures are repeated for different choices of the regularization parameters $\{a,\gamma_{1},\gamma_{2}\}$ and results are reported in fig. 5. Specifically, fig. 5a is a colour plot showing the performance of the network in the plane $(sD_{KL},C_{v}(1))$, where each point is coloured according to the slope $\omega$. The three lines correspond to the three regularizations: the unlearning line is generally closer to the origin than the $L_{1}$ and $L_{2}$ ones, suggesting a better compromise between accuracy and robustness. All the lines converge to the ground-truth when $sD_{KL}=0$, corresponding to absence of regularization ($a=1$ or $\gamma_{1}=0$ or $\gamma_{2}=0$).
@@ IV.4 SK model with early-stopping
Figure 7: Comparison at early-stopping of the performance of the unlearning, $L_{1}$ and $L_{2}$ regularization schemes. Each point has two coordinates: the symmetric Kullback-Leibler divergence $sD_{KL}$ and specific heat at unitary temperature $C_{v}(1)$. The colorbar relates each point to the standard deviation of the couplings $\sigma_{J}$. Arrows display the increase in the regularization rates of the algorithms. The dashed line indicates the specific heat of the original SK model that generated the data. Choice of the parameters: $N=18$, $\beta_{d}=0.4$, $a\in[0,0.63]$ in the U upper branch and $a\in[1.9,3]$ in the U lower branch, $\gamma_{1}\in[0.5,0.4]$, $\gamma_{2}\in[0.3,4]$, $\lambda=0.01$, $\lambda_{1}=\lambda_{2}=0.001$.
@@
IV.5 SK model with $N\gg 1$
In this section, we test the method on an instance with larger number $N$ of variables, showing a good efficiency of the Unlearning method, both run until convergence and early-stopped.
A number $M=10^{4}$ of data-points are sampled from a SK model with $N=200$ spins in equilibrium at a temperature $\beta_{d}=0.4$ (i.e. in the paramagnetic phase). The network is then initialized in Hebbian fashion, i.e. $\bm{J}^{(0)}=\bm{c}_{d}$ and $\bm{h}^{(0)}=\bm{0}$. The model is trained via Unlearning, $L_{1}$ and $L_{2}$ regularization methods and the specific heat $C_{v}(1)$, the slope $\omega(\bm{c}_{1},\bm{c}_{d})$ and the Pearson coefficient $\rho(\bm{c}_{1},\bm{c}_{d})$ are measured through MCMC sampling at unitary temeperature. Two different training prescriptions have been adopted for the study: the algorithm is executed until convergence of equations (27), (28); training is early-stopped as soon as $\omega\simeq 1$.
Figure 9: Comparison of the performance of the unlearning with $L_{1}$ and $L_{2}$ regularization methods after training with $M=10^{4}$ configurations of a SK model with $N=200$ spins and inverse temperature $\beta_{d}=0.4$. (a) The Pearson coefficient $\rho(c_{1},c_{d})$ is plotted as a function of $\omega(c_{1},c_{d})$ for different values of specific heat at unitary temperature $C_{v}(1)$. The dashed lines represent the best (i.e. unitary) values for $\rho$ and $\omega$, that meet in the non-regularized performance. Choice of the parameters: $\lambda=\lambda_{1}=\lambda_{2}=10^{-2}$ $a\in[0.45,1]$, $\gamma_{1}\in[0,0.09]$, $\gamma_{2}\in[0,4.5]$. (b) $C_{v}(1)$ as a function of $\rho(c_{1},c_{d})$ when the algorithm is early stopped at $\omega\simeq 1$. Measures are derived from one training instance, since they resulted stable enough over different repetitions. Choice of the parameters: $\lambda=\lambda_{1}=\lambda_{2}=10^{-3}$ $a\in[0.45,0.9]$, $\gamma_{1}\in[0.01,0.18]$, $\gamma_{2}\in[0.5,9]$.
Figure 9 depicts the values of $\rho$ as a function of $\omega$ and $C_{v}(1)$ as achieved through the three regularization methods for various values of the control parameters $a,\gamma_{1},\gamma_{2}$. At convergence (see fig. 9a), the Unlearning algorithm appears to approach a better accuracy performance (i.e. in terms of both $\rho$ and $\omega$) together with a lower value of the specific heat. Even after early-stopping (see fig. 9b) the same behaviour emerges: there is a range in $a$ where, given the same degree of accuracy, the model achieves a smaller specific heat with respect of the other regularization methods. We also stress that, by contrast with the rest of our analysis, this experiment provides for a specific heat $C_{v}(1)$ obtained at early-stopping which is lower than the one reached at convergence. This aspect might be due to the Hebbian initialization of the parameters in a regime which is very far from the perfect-sampling one that we considered before.
It should be noted that the standard deviation $\sigma_{J}$ of the couplings is chosen to have the same order of magnitude, across all three regularization methods. Both in the converged and early-stopped scenario, the Unlearning algorithm is superior to the $L_{p}$ regularizations, confirming the results obtained at small $N$ (i.e. the perfect sampling framework). Moreover, the early-stopping technique used in this work, which relies on the Hebbian initialization of the network, results handily applicable for larger dimensions due to its reduced computational cost, hence providing an effective training recipe in terms of both accuracy and robustness.